# Compressing neural network by tensor network with exponentially fewer variational parameters

### v2

@@ Abstract
Neural network (NN) designed for challenging machine learning tasks is in general a highly nonlinear mapping that contains massive variational parameters. High complexity of NN, if unbounded or unconstrained, might unpredictably cause severe issues including over-fitting, loss of generalization power, and unbearable cost of hardware. In this work, we propose a general compression scheme that significantly reduces the variational parameters of NN by encoding them to deep automatically-differentiable tensor network (ADTN) that contains exponentially-fewer free parameters. Superior compression performance of our scheme is demonstrated on several widely-recognized NN’s (FC-2, LeNet-5, AlextNet, ZFNet and VGG-16) and datasets (MNIST, CIFAR-10 and CIFAR-100). For instance, we compress two linear layers in VGG-16 with approximately $10^{7}$ parameters to two ADTN’s with just 424 parameters, where the testing accuracy on CIFAR-10 is improved from $90.17\%$ to $91.74\%$. Our work suggests TN as an exceptionally efficient mathematical structure for representing the variational parameters of NN’s, which exhibits superior compressibility over the commonly-used matrices and multi-way arrays.
@@ I Introduction
Neural network (NN) [1] has gained astounding success in a wide range of fields including computer vision, natural language processing, and most recently scientific investigations in, e.g., applied mathematics(e.g.,  [2, 3]) and physics(e.g., [4, 5, 6, 7, 8, 9, 10]). To improve the performance on handling complicated realistic tasks, the amount of variational parameters in NN rapidly increases from millions to trillions (e.g., Chat-GPT with 175 billion parameters [11]). Such a paradigm with the utilization of super large-scale NN’s brought us several severe challenges. Though the representation ability of NN should be enhanced by increasing the complexity, over-fitting might occur and the generalization ability of NN might unpredictably be harmed. The increasing complexity also brings unbearable cost of hardware in academic investigations and practical applications.
@@
The variational parameters in NN are usually stored as matrices or higher-order tensors (also called multi-linear arrays). Previous results show that generalization ability can be improved by suppressing or “refining” the degrees of freedom in such tensors by, e.g., network pruning [12, 13, 14], knowledge distillation [15, 16], weight sharing [17, 18], tensor decompositions [19, 20], etc. In recent years, shallow tensor networks (TN’s) including matrix product state (MPS [21, 22], also known as tensor-train form [23]) and matrix product operator (MPO [24]) were applied to represent the tensors in neural networks, and achieved remarkable compression rates [25, 26, 27, 28, 29]. These imply TN as a superior mathematical structure over the simple multi-linear arrays for representing the variational parameters of NN’s, which yet remains an open issue.
@@
In this work, we propose to encode the variational parameters of the considered NN layer(s) into the contraction of TN [30] that contains exponentially fewer parameters. See the illustration of the main procedures in Fig. 1. For instance, the number of parameters of the TN encoding $2^{Q}$ parameters of NN scale just linearly as $O(MQ)$, with $M\sim O(1)$ the number of TN layers. Since the contraction process is differentiable, automatic differentiation technique [31, 32] is utilized to optimize the tensors of the TN to reach the optimal accuracy after compression. Thus, we dub our scheme as automatically differentiable tensor network (ADTN).
@@
We demonstrate the compression performance of ADTN on various well-known NN’s (FC-2, LeNet-5 [33], AlexNet, ZFnet, and VGG-16 [34]) and datasets (MNIST [35], CIFAR-10 [36] and CIFAR-100). Note the details of NN’s can be found in Appendix A. For instance, in the experiment of compressing two fully connected layers of VGG-16, we compressed about $10^{7}$ parameters to two ADTN’s with just 424, where the accuracy of the testing set of CIFAR-10 is improved from $90.17\%$ to $91.74\%$. Our work implies a general way to reduce the excessive complexity of NN’s by encoding the variational parameter into TN as a more compact and efficient mathematical form.
II Method
A NN can be formally written as a mapping
@@
Denoting the parameters to be compressed as tensor $\boldsymbol{T}$, our idea is to represent $\boldsymbol{T}$ as the contraction of ADTN with multiple layers. The structure of ADTN is flexible. Here, we choose the commonly-used “brick-wall” structure; see an example of its diagrammatic representation in Fig. 1 (b). In simple words about the diagrammatic representation of TN, each block represents a tensor, and the bonds connected to a block represent the indexes of the corresponding tensor. Shared bonds represent the dumb indexes that should be summed up (contracted). The vertical black lines represent the activation function (e.g., ReLU). More details about TN contractions and the diagrammatic representation can be found in Appendix A or Ref. [30].
@@
The exemplified ADTN is formed by several $(d\times d\times d\times d)$ tensors $\{\boldsymbol{A}^{[k]}\}$ ($k=0,\ldots,K-1$ with $K$ their total number), which form the variational parameters of ADTN. On the left boundary of ADTN, we put several two-dimensional constant vectors $\boldsymbol{v}$ for the convenience in formulating and analyzing. We may take $\boldsymbol{v}=[1,0]$ for instance.
@@
There are two main stages to obtain the tensors in ADTN $\{\boldsymbol{A}^{[k]}\}$. The first is pre-training by minimizing the distance between $\boldsymbol{T}$ and $\boldsymbol{\mathcal{T}}$. We choose the loss function as Euclidean distance $L^{\text{E}}=|\boldsymbol{\mathcal{T}}-\boldsymbol{T}|$. After $L^{\text{E}}$ converges, we move to the second stage to minimize the loss function for implementing the machine learning task. One may choose the same loss function as the one used for training the NN, such as the cross entropy from the outputs of NN and the ground-truth classifications of the training samples. The feed-forward process of getting the output of the NN is the same as that for training the NN, except that the parameters ($\boldsymbol{T}$) of the target NN layer are replaced by the tensor ($\boldsymbol{\mathcal{T}}$) obtained by contracting the ADTN. Generally speaking, the first stage is to enhance the stability by finding a proper initialization of ADTN, and the second stage further improves the performance with an end-to-end optimization. The mapping given by the NN in the second stage can be written as
@@
The number of compressed NN layers and that of ADTN’s are flexible. The mapping of the compressed NN can be written as $\boldsymbol{y}=f(\boldsymbol{x};\{\boldsymbol{A}^{[k]}\},\{\boldsymbol{B}^{[k]}\},\cdots,\boldsymbol{T}^{\text{res}})$, with $\{\boldsymbol{A}^{[k]}\},\{\boldsymbol{B}^{[k]}\},\cdots$ denoting the tensors of multiple ADTN’s. This allows more convenient treatment on the cases where the number of the parameters cannot be written as $d^{Q}$. For instance, $(3\times 2^{10})$ parameters in a NN layer can be compressed to a single ADTN with totally $Q=11$ unshared indexes, where 10 of them are two-dimensional and 1 is three-dimensional. It can also be divided into two parts with $2^{11}$ and $2^{10}$ parameters, respectively, which are then compressed to two ADTN’s with $Q=11$ and $10$, respectively. All unshared indexes in these two ADTN’s can be two-dimensional.
III Compression performance
@@
Table 1: ADTN’s performance on compressing the linear layers (green area) and convolutional layers (oriange area) of FC-2, LeNet-5, Alex Net, ZF Net, and VGG-16 for MNIST, CIFAR-10, and CIFAR-100 datasets. We show the number of parameters in the compressed NN layer(s) $P$, that of ADTN(s) $\mathcal{P}$, compression ratio $\rho=\mathcal{P}/P$, the number of ADTN(s) $N$, and testing accuracies before compression $\eta_{\text{NN}}$ and after compression $\eta$. The hyper-parameters of these NN’s can be found in Appendix B.
@@
In Table 1, the green area shows the performance of ADTN with $M=1$ to compress linear layers in various NN’s. High compression ratios $\rho=\mathcal{P}/P$ (with $P$ and $\mathcal{P}$ the numbers of parameters in the compressed NN layer(s) and in the ADTN(s), respectively) are achieved. Our first example is FC-2 formed by two linear layers, whose sizes are $(784\times 256)$ and $(256\times 10)$, respectively. We slice out $2^{17}$ parameters from the first linear layer and compress them in an ADTN with just 160 parameters. The testing accuracy on the MNIST dataset, which consisting of 60000 hand-written images of size $(28\times 28)$ [35], remains almost unchanged after compression. In more complex convolutional NN’s (LeNet-5, AlexNet, and ZFNet), we chose to compress all linear layers except for the output layer. This is because the number of parameters in the output layer is much smaller compared to other layers. For different datasets (MNIST, CIFAR-10, and CIFAR-100), we just accordingly alter the dimensions for the input and output of each NN, and keep other hyper-parameters unchanged. Note CIFAR-10 and CIFAR-100 contain totally 80 million everyday-life RGB images of size $(32\times 32)$ [36]. The ADTN’s with just hundreds of parameters manage to compress the linear layers whose numbers of parameters vary from $O(10^{3})$ to $O(10^{7})$, while the testing accuracy is not harmed by compression. Remarkably in compressing VGG-16, we use two ADTN’s with in total 424 parameters to compress two linear layers with approximately $10^{7}$ parameters (whose dimensions are $(512\times 4096)$ and $(4096\times 4096)$, respectively), and meanwhile are able to increase the testing accuracy from $90.17\%$ to $91.74\%$.
@@
The orange area in Table 1 shows excellent performance of ADTN to compress convolutional layers. For example, we compress the two largest convolutional layers in AlexNet and ZFNet (with about $O(10^{6})$ and $O(10^{5})$ parameters) that account for more than $80\%$ of the total parameters in the corresponding NN. These parameters are compressed to the ADTN’s consisting of just 324 parameters. For VGG-16, we choose to compress the five largest layers from the $13$ convolutional layers. Each convolutional layer used two ADTN’s to compress. The testing accuracy is improved from about 1 percentage.
@@
We shall emphasize that many settings of our ADTN compression scheme are flexible, such as the choices of compressed NN layers, the number of ADTN’s, and the hyper-parameters of each ADTN. The demonstrated results in this paper are obtained after slightly adjusting these quantities. Higher compression performance can generally be reached by further adjusting them.
@@
One can consider to simultaneously compress multiple types of NN layers. In Table 2, we show the performance of ADTN for simultaneously compressing linear and convolutional layers of VGG-16. We use a reduced version of VGG-16 (with smaller sizes of the NN layers and without data augmentation) for demonstration, with its testing accuracy $\eta_{\text{NN}}=81.14\%$. The red bars in the first column indicate the compressed layers, and their lengths illustrate the numbers of parameters. Here, each NN layer is compressed to one ADTN. Remarkable compression rations ($\rho$) are achieved.
@@
Table 2: (Color online) The performance of ADTN scheme with different numbers of tensor layers ($M$) to compress both the linear and convolutional layers of a reduced version of VGG-16 (with its testing accuracy $\eta_{\text{NN}}=81.14\%$). In the first column, the red bars indicate the compressed layers and their lengths illustrate the numbers of parameters. The second and third columns show the training and testing accuracies ($\tilde{\eta}$ and $\eta$, respectively). The last column gives the total compression ratio $\rho_{\text{tot}}$ [Eq. (5)].

### v3

# Compressing Neural Networks Using Tensor Networks with Exponentially Fewer Variational Parameters

@@ Abstract
Neural network (NN) designed for challenging machine learning tasks is in general a highly nonlinear mapping that contains massive variational parameters. High complexity of NN, if unbounded or unconstrained, might unpredictably cause severe issues including overfitting, loss of generalization power, and unbearable cost of hardware. In this work, we propose a general compression scheme that significantly reduces the variational parameters of NN’s, despite of their specific types (linear, convolutional, etc), by encoding them to deep automatically differentiable tensor network (ADTN) that contains exponentially-fewer free parameters. Superior compression performance of our scheme is demonstrated on several widely-recognized NN’s (FC-2, LeNet-5, AlextNet, ZFNet and VGG-16) and datasets (MNIST, CIFAR-10 and CIFAR-100). For instance, we compress two linear layers in VGG-16 with approximately $10^{7}$ parameters to two ADTN’s with just 424 parameters, improving the testing accuracy on CIFAR-10 from $90.17\%$ to $91.74\%$. We argue that the deep structure of ADTN is an essential reason for the remarkable compression performance of ADTN, compared to existing compression schemes that are mainly based on tensor decompositions/factorization and shallow tensor networks. Our work suggests deep TN as an exceptionally efficient mathematical structure for representing the variational parameters of NN’s, which exhibits superior compressibility over the commonly-used matrices and multi-way arrays.
@@ I Introduction
Neural network (NN) [1] has achieved remarkable success across a wide range of fields including computer vision, natural language processing, and most recently scientific investigations in, e.g., applied mathematics (e.g.,  [2, 3]) and physics (e.g.,  [4, 5, 6, 7, 8, 9, 10]). To enhance performance in handling complex real-world tasks, such as human interaction (e.g., [11]) and robot control (e.g.,[12, 13, 14]), the amount of variational parameters in NN rapidly increased from millions to trillions (e.g., GPT-3, with 175 billion parameters [15]). Such a paradigm with the utilization of super large-scale NN’s brought us several severe challenges. Though the representation ability of NN should be enhanced by increasing complexity, Roverfitting might occur and the generalization ability of NN might unpredictably be harmed. The increasing complexity also brings unbearable hardware costs in academic investigations and practical applications.
@@
The variational parameters in NN are usually stored as matrices or higher-order tensors (also called multilinear arrays). Previous results show that generalization ability can be improved by suppressing or “refining” the degrees of freedom in such tensors by, e.g., network pruning [16, 17, 18], knowledge distillation [19, 20], weight sharing [21, 22], tensor Rdecompositions and factorizations [23, 24], etc.
Among the inspiring progresses made in recent years, a simple and efficient iterative approach known as MUSCO[25] alternates the low-rank factorization with smart rank selection and fine-tuning. Mixed tensor decomposition (MTD) [26] attempts to combine Tucker decomposition and canonical polyadic decomposition (CPD) to for better performance. Low-rank adaptation (LoRA) [27] freezes all pre-trained parameters and inserts a trainable pair of matrices (acting as a low-rank decomposition of a full matrix) additively into each layer of the Transformer architecture.
Meanwhile, tensor network (TN) demonstrates its significant prospects in model compression. Matrix product state (MPS) [28, 29], which is also known as tensor-train (TT) form [30], and matrix product operator (MPO) [31], were applied to compress the tensors in NN’s, and achieved remarkable compression ratios [32, 33, 34, 35, 36]. Nonlinear TT [37] was introduced by containing nonlinear activation functions embedded in the sequenced contractions and convolutions on the top of the normal TT decomposition. These imply TN as a superior mathematical structure over simple multilinear arrays for representing the variational parameters of NN’s, which yet remains an open issue. However, these works mainly concern to use the shallow TN’s. The utilization of deep TN’s, which were revealed to be essentially different and more powerful from the research in quantum physics and machine learning, has not been investigated yet for model compression.
@@
In this work, we propose to encode the variational parameters of NN layer(s) into the contraction of TN [38] that contains exponentially fewer parameters. See the illustration of the main procedures in Fig. 1. For instance, the number of parameters of the TN encoding $2^{Q}$ parameters of NN scale just linearly as $O(MQ)$, with $M\sim O(1)$ the number of TN layers. Since the contraction process is differentiable, automatic differentiation technique [39, 40] is utilized to optimize the tensors of the TN to reach the optimal accuracy after compression. Thus, we dub our scheme as automatically differentiable tensor network (ADTN).
ADTN exhibits fundamental differences from previous compression methods. Unlike traditional approaches that rely on explicit decomposition or factorization procedures or the insertion of matrices, ADTN directly encodes NN parameters into a deep TN and updates them through automatic differentiation techniques. This distinguishes ADTN from the TT-based methods and LoRA. By utilizing deep TN’s, ADTN leverages a distinct representational power that surpasses the capabilities of shallow TN’s used in the existing tensor decompositions, factorizations, and MPS/MPO methods. We will give more discussions later in Sec. 4.
@@
We demonstrate the compression performance of ADTN on various well-known NN’s (FC-2, LeNet-5 [41], AlexNet, ZFnet, and VGG-16 [42]) and datasets (MNIST [43], CIFAR-10 and CIFAR-100 [44]). Note that the details of NN’s can be found in Appendix A. For instance, in the experiment of compressing two fully connected layers of VGG-16, we compressed about $10^{7}$ parameters to two ADTN’s with just 424 parameters, where the testing accuracy of CIFAR-10 improved from $90.17\%$ to $91.74\%$. Our work implies a general way to reduce the excessive complexity of NN’s by encoding the variational parameter into TN’s, which serve as a more compact and efficient mathematical form.
II Methods
II.1 Encoding variational parameters of NN’s to ADTN’s
An NN can be formally written as a mapping
@@ II.1 Encoding variational parameters of NN’s to ADTN’s
Denoting the parameters to be compressed as tensor $\boldsymbol{T}$, our idea is to represent $\boldsymbol{T}$ as the contraction of ADTN with multiple layers. The structure of ADTN is flexible. Here, we choose the commonly used “brick-wall” structure; see an example of its diagrammatic representation in Fig. 1 (b). In simple words about the diagrammatic representation of TN, each block represents a tensor, and the bonds connected to a block represent the indexes of the corresponding tensor. Shared bonds represent the dumb indexes that should be summed up (contracted). The vertical black lines represent the activation function (e.g., ReLU). More details about TN contractions and the diagrammatic representation can be found in Appendix A or Ref. [38].
@@
The exemplified ADTN is formed by several $(d\times d\times d\times d)$ tensors $\{\boldsymbol{A}^{[k]}\}$ ($k=1,\ldots,K$ with $K$ as their total number), which form the variational parameters of ADTN. On the left boundary of ADTN, we put several two-dimensional constant vectors $\boldsymbol{v}$ for the convenience in formulating and analyzing. We may take $\boldsymbol{v}=[1,0]$ for instance.
@@
[…]
By encoding $2^{Q}$ parameters, the total number of variational parameters in the brick-wall ADTN depicted in Fig. 1 (b) scales as
@@
II.2 Optimization of ADTN
There are two main stages to obtain the tensors $\{\boldsymbol{A}^{[k]}\}$ in an ADTN. The first is pre-training by minimizing the distance between $\boldsymbol{T}$ in NN and $\boldsymbol{\mathcal{T}}$ obtained from ADTN. We choose the loss function as Euclidean distance $L^{\text{E}}=|\boldsymbol{\mathcal{T}}-\boldsymbol{T}|$, as the parameter values of the neural network lack specific meaning. For more on the impact of different pretraining strategies, see Appendex A.
After $L^{\text{E}}$ converges, we move to the second stage to minimize the loss function for implementing the machine learning task. One may choose the same loss function as the one used for training the NN, such as the cross entropy from the outputs of NN and the ground-truth classifications of the training samples. The feed-forward process of getting the output of the NN is the same as that for training the NN, except that the parameters ($\boldsymbol{T}$) of the target NN layer are replaced by the tensor ($\boldsymbol{\mathcal{T}}$) obtained by contracting the ADTN.
Generally speaking, the first stage is to enhance the stability by finding a proper initialization of ADTN, and the second stage further improves the performance with an end-to-end optimization. The mapping given by the NN in the second stage can be written as
@@ II.2 Optimization of ADTN
The number of compressed NN layers, the number of ADTN’s, and the number of TN layers within the ADTN’s are all flexible. The mapping of the compressed NN can be expressed as $\boldsymbol{y}=f(\boldsymbol{x};\{\boldsymbol{A}^{[k]}\},\{\boldsymbol{B}^{[k]}\},\allowbreak\cdots,\boldsymbol{T}^{\text{res}})$, with $\{\boldsymbol{A}^{[k]}\},\{\boldsymbol{B}^{[k]}\},\cdots$ denoting the tensors of multiple ADTN’s. This flexibility facilitates handling cases where the number of parameters cannot be expressed as $d^{Q}$. For instance, $(3\times 2^{10})$ parameters in a NN layer can be compressed into a single ADTN with $Q=11$ unshared indexes, where 10 are two-dimensional and one is three-dimensional. Alternatively, these parameters can also be divided into two parts with $2^{11}$ and $2^{10}$ parameters, respectively, which are then compressed to two ADTN’s with $Q=11$ and $10$, respectively. All unshared indexes in these two ADTN’s can be two-dimensional.
Let us provide a concrete example to explain the choice of $Q$ and the number of ADTNs. We may consider the compression of a $(600\times 600)$ weight matrix (denoted as $\boldsymbol{W}$) in a linear layer. One can consider compressing a $(512\times 512)$ part of $\boldsymbol{W}$, which contains $2621442^{18}=$ 262144 parameters (with $Q=18$). This leaves $360000-262144=97856$ parameters uncompressed, where one can optionally introduce another ADTN of $Q=16$ to compress $655362^{16}=$ 65536 parameters. Following this approach, an appropriate number of ADTNs can be introduced to ensure that most (or a desired portion) of a given parameter tensor is compressed into ADTNs. Another way to compress the entire $\boldsymbol{W}$ with a single ADTN is to choose $Q=19$, with $3600002^{19}=524288>$ 360000. One could add zeros to $\boldsymbol{W}$ to increase the number of parameters to $2^{19}$. But notably, the number of parameters in the $Q=19$ ADTN will still be much lower than that of $\boldsymbol{W}$.
ADTN distinguishes itself from conventional compression techniques such as network distillation, model quantization [45], and low-rank decompositions. Network distillation [16, 17, 18] transfers knowledge from a larger model to a smaller one, while model quantization reduces numerical precision to minimize memory and computational costs. Low-rank decompositions  [23, 24] approximate high-dimensional data by representing it with lower-rank approximations, thereby reducing parameter complexity. In contrast, ADTN directly encodes the parameter tensors into deep TN’s. The ADTN method is independent of network distillation and model quantization, though these techniques can be integrated into ADTN’s pre-training and fine-tuning stages to further enhance compression. ADTN can be considered a lower-rank structure with a deep architecture, distinguishing it from the relatively shallow structures typically produced by conventional low-rank decompositions.
III Results and Discussions
III.1 Compression Performance
For all the ADTN’s used in this work, we take the brick-wall structure, where all tensors are $(2\times 2\times 2\times 2)$. The activation functions are taken as ReLU. The value of $Q$ is taken according to the size of the compressed tensor. We only vary the number of ADTN’s ($N$) and the number of TN layers ($M$) in each ADTN in the compression of each NN layer. The specific values of $N$ and $M$ are given in the corresponding texts describing the results.
Dataset NN Model #Parameters Compression ratio $\rho=\mathcal{P}/P$ADTN $N$ Testing accuracy
Compressed
NN layer(s) $P$
ADTN(s)
$\mathcal{P}$
Baseline
$\eta_{NN}$
$\eta$ $\eta/\eta_{NN}$
Compress linear layers(s) MNIST FC-2 $131072$ $160$ $1.2\times 10^{-3}$ $1$ $97.94\%$ $97.81\%$ $99.69\%$
LeNet-5 $8192$ $120$ $1.5\times 10^{-2}$ $1$ $99.15\%$ $99.48\%$ $100.36\%$
CIFAR-10 LeNet-5 $163840$ $300$ $1.8\times 10^{-3}$ $2$ $74.39\%$ $75.27\%$ $101.18\%$
AlexNet $5242880$ $388$ $7.4\times 10^{-5}$ $2$ $81.13\%$ $81.31\%$ $100.22\%$
ZFNet $6291456$ $404$ $6.4\times 10^{-5}$ $2$ $80.59\%$ $81.03\%$ $100.66\%$
VGG-16 $18874368$ $424$ $2.2\times 10^{-5}$ $2$ $90.17\%$ $91.74\%$ $101.74\%$
CIFAR-100 LeNet-5 $163840$ $300$ $1.8\times 10^{-3}$ $2$ $41.30\%$ $42.60\%$ $103.15\%$
AlexNet $5242880$ $388$ $7.4\times 10^{-5}$ $2$ $49.80\%$ $49.19\%$ $98.78\%$
ZFNet $6291456$ $404$ $6.4\times 10^{-5}$ $2$ $50.19\%$ $50.36\%$ $100.34\%$
Compress conv layers(s) MNIST LeNet-5 $32768$ $140$ $4.2\times 10^{-3}$ $1$ $99.15\%$ $99.52\%$ $100.37\%$
CIFAR-10 LeNet-5 $8192$ $120$ $1.5\times 10^{-2}$ $1$ $74.39\%$ $75.16\%$ $101.04\%$
AlexNet $1572864$ $364$ $2.9\times 10^{-4}$ $2$ $81.13\%$ $81.09\%$ $99.95\%$
ZFNet $393216$ $324$ $8.2\times 10^{-4}$ $2$ $80.59\%$ $81.63\%$ $101.29\%$
VGG-16 $11796480$ $1820$ $1.5\times 10^{-4}$ $10$ $90.17\%$ $91.09\%$ $101.02\%$
CIFAR-100 LeNet-5 $8192$ $120$ $1.5\times 10^{-2}$ $1$ $41.30\%$ $42.54\%$ $103.00\%$
AlexNet $1572864$ $364$ $2.9\times 10^{-4}$ $2$ $49.80\%$ $50.15\%$ $100.70\%$
ZFNet $393216$ $324$ $8.2\times 10^{-4}$ $2$ $50.19\%$ $50.88\%$ $101.37\%$
@@ III.1 Compression Performance
Table 1: ADTN’s performance on compressing the linear layers (green area) and convolutional layers (orange area) of FC-2, LeNet-5, Alex Net, ZF Net, and VGG-16 for MNIST, CIFAR-10, and CIFAR-100 datasets. We show the number of parameters in the compressed NN layer(s) $P$, that of ADTN(s) $\mathcal{P}$, compression ratio $\rho=\mathcal{P}/P$, the number of ADTN(s) $N$, and testing accuracies before compression $\eta_{\text{NN}}$ and after compression $\eta$. The hyper-parameters of these NN’s can be found in Appendix B.
In Table 1, the green area shows the performance of ADTN with $M=1$ for compressing linear layers in various NN’s. High compression ratios $\rho=\mathcal{P}/P$ (with $P$ and $\mathcal{P}$ the number of compressed parameters in the NN layer(s) and that of the ADTN(s), respectively) are achieved.
Our first example is FC-2 formed by two linear layers, whose sizes are $(784\times 256)$ and $(256\times 10)$, respectively. We slice out $2^{17}$ parameters from the first linear layer and compress them in an ADTN with just 160 parameters. The testing accuracy on the MNIST dataset, which consisting of 60000 hand-written images of size $(28\times 28)$ [43], remains almost unchanged after compression.
@@
In more complicated convolutional NN’s (LeNet-5, AlexNet, and ZFNet), we chose to compress all linear layers except for the output layer. This is because the number of parameters in the output layer is much smaller compared to other layers. For different datasets (MNIST, CIFAR-10, and CIFAR-100), we just accordingly alter the dimensions for the input and output of each NN, and keep other hyper-parameters unchanged. Note CIFAR-10 and CIFAR-100 contain totally 80 million everyday-life RGB images of size $(32\times 32)$ [44].
The ADTN’s with just hundreds of parameters manage to compress the linear layers whose numbers of parameters vary from $O(10^{3})$ to $O(10^{7})$, while the testing accuracy is not harmed by compression. Remarkably in compressing VGG-16, we use two ADTN’s with in total 424 parameters to compress two linear layers with approximately $10^{7}$ parameters (whose dimensions are $(512\times 4096)$ and $(4096\times 4096)$, respectively), and meanwhile are able to increase the testing accuracy from $90.17\%$ (baseline) to $91.74\%$.
@@
The orange area in Table 1 shows the excellent performance of ADTN to compress convolutional layers. For example, we compress the two largest convolutional layers in AlexNet and ZFNet (with about $O(10^{6})$ and $O(10^{5})$ parameters) that account for more than $80\%$ of the total parameters in the corresponding NN. These parameters are compressed to ADTN’s consisting of just 324 parameters. For VGG-16, we choose to compress the five largest layers from the $13$ convolutional layers. Each convolutional layer used two ADTN’s to compress. The testing accuracy is improved for about 1 percentage.
@@
We shall emphasize that many settings of our ADTN compression scheme are flexible, such as the choices of compressed NN layers, the number of ADTN’s, and the hyper-parameters of each ADTN. The demonstrated results in this paper are obtained after slightly adjusting these quantities. Higher compression performance can generally be reached by adjusting them more carefully.
@@
One can consider to simultaneously compress multiple types of NN layers. In Table 2, we show the performance of ADTN for simultaneously compressing linear and convolutional layers of VGG-16. We use a reduced version of VGG-16 (with smaller sizes of the NN layers and without data augmentation) for demonstration, with its baseline testing accuracy $\eta_{\text{NN}}=81.14\%$. In the figure at the left side of Table  2, the red bars indicate the compressed layers, and their lengths illustrate the numbers of parameters, while the green bars represent the uncompressed layers. Here, each NN layer is compressed to one ADTN. Remarkable compression ratios ($\rho$) are achieved.
Compressed
layers(red)
M $\tilde{\eta}$ $\eta$ $\eta/\eta_{NN}$ $\rho$ $\rho_{tot}$
(a) Conv 1 $99.87\%$ $84.27\%$ $103.86\%$ $5.0\times 10^{-5}$ $0.435$
3 $100\%$ $85.52\%$ $105.40\%$ $2.1\times 10^{-4}$
5 $100\%$ $85.71\%$ $105.63\%$ $3.7\times 10^{-4}$
(b) Conv 1 $99.99\%$ $84.36\%$ $103.97\%$ $6.0\times 10^{-5}$ $0.279$
3 $100\%$ $85.19\%$ $105.00\%$ $2.5\times 10^{-4}$
5 $100\%$ $85.68\%$ $105.60\%$ $4.4\times 10^{-4}$
(c) Conv 1 $98.83\%$ $78.53\%$ $99.78\%$ $8.0\times 10^{-5}$ $0.135$
3 $99.93\%$ $82.44\%$ $101.60\%$ $3.3\times 10^{-4}$
5 $100\%$ $82.79\%$ $102.03\%$ $5.9\times 10^{-4}$
(d) Conv +Linear 1 $96.35\%$ $65.93\%$ $81.25\%$ $1.0\times 10^{-4}$ $0.060$
3 $99.94\%$ $77.73\%$ $95.78\%$ $4.2\times 10^{-4}$
5 $100\%$ $79.15\%$ $97.55\%$ $7.4\times 10^{-4}$
7 $100\%$ $80.61\%$ $99.35\%$ $1.1\times 10^{-3}$
(e) Conv +Linear 3 $75.78\%$ $72.13\%$ $88.90\%$ $1.0\times 10^{-4}$ $0.008$
5 $78.13\%$ $74.24\%$ $91.50\%$ $4.2\times 10^{-4}$
7 $81.54\%$ $77.61\%$ $95.65\%$ $7.4\times 10^{-4}$
9 $81.95\%$ $78.59\%$ $96.86\%$ $1.1\times 10^{-3}$
@@
Figure 2: (Color online) The performance of ADTN scheme with different numbers of tensor layers ($M$) to compress both the linear and convolutional layers of a reduced version of VGG-16 (with the baseline testing accuracy $\eta_{\text{NN}}=81.14\%$). In the figure at the left side, the red bars indicate the compressed layers, and their lengths illustrate the numbers of parameters, while the green bars represent the uncompressed layers. In the table at the right side, the three and four columns show the training and testing accuracies ($\tilde{\eta}$ and $\eta$, respectively). The last column gives the total compression ratio $\rho_{\text{tot}}$ [Eq. (5)].
@@
The approximation holds since in general we have #$\#(\text{ADTN})\ll\#(\boldsymbol{T}^{\text{res}})\ll\#(\text{NN})$. Therefore, $\rho_{\text{tot}}$ can be significantly improved by simply compressing more NN layers to ADTN’s.